an angles complement is 24. how many degrees does the angles supplement measure?
Found 2 solutions by dabanfield, richwmiller:
Answer by dabanfield(803)

an angles complement is 24. how many degrees does the angles supplement measure?
Let x be the complement of 24.
Then x + 24 = 90
So we have x = 90-24 = 56
Let y be the supplement of 56.
Then y + 56 = 180
y = 180 - 56 = 124
The answer is 124.

Answer by richwmiller(17219)

90-24=66 not 56
so 180-66= 114 supplement